# Redistribution Through

### v2

@@ 2 Model
The value $w(\theta,p)$ thus represents the individual contribution to the aggregate welfare when a buyer of type $\theta$ is assigned to a segment where the seller charges price $p$.
We restrict attention to welfare functions satisfying three conditions. First, the value from assigning a buyer to a segment where he does not consume is zero, and the value from assigning a buyer to a segment where he consumes is non-negative. That is, $w(\theta,p)=0$ if $\theta<p$, and $w(\theta,p)\geq 0$ if $\theta\geq p$. Second, assigning any buyer to a segment where the seller charges a lower price is always welfare-improving. That is, $p\mapsto w(\theta_{k},p)$ is a non-increasing function on $\{\theta_{1},\dots,\theta_{k}\}$ for all $k\in\{1,\dots,K\}$. The two preceding conditions capture the fact that social welfare aligns with consumer surplus. [^16] It is straightforward to verify that the buyer-utilitarian welfare function, given by $w(\theta,p)=(\theta-p)\mathds{1}_{\theta\geq p}$ for all $(\theta,p)\in\Theta\times P$, satisfies these two minimal conditions. The third condition captures the designer’s redistributive concerns. We require that any welfare function is supermodular conditional on buyers consuming. That is, for any $(\theta,p),(\theta^{\prime},p^{\prime})\in\Theta\times P$ such that $\theta^{\prime}>\theta\geq p^{\prime}>p$,
@@
Condition R ensures that any price discount increases welfare more when it is granted to low-type buyers rather than high-types.
We say that a welfare function is redistributive if it satisfies the three conditions above and denote as $\mathcal{R}$ the class of those welfare functions. Furthermore, we say that $w$ is strictly redistributive if it satisfies condition R strictly on all $\Theta\times P$ and is such that $p\mapsto w(\theta_{k},p)$ is strictly decreasing on $\{\theta_{1},\dots,\theta_{k}\}$ for all $k\in\{1,\dots,K\}$. We denote the class of strictly redistributive welfare functions as $\overline{\mathcal{R}}$.
In Examples 1 and 2 below, we highlight that the class of redistributive welfare functions encompasses two common specifications.
@@
belongs to $\mathcal{R}$ whenever $\theta\mapsto\lambda(\theta)$ is non-increasing, and belongs to $\overline{\mathcal{R}}$ whenever $\theta\mapsto\lambda(\theta)$ is strictly decreasing. The class of welfare functions given by Equation 1 for some decreasing function $\lambda$ correspond to the family of weighted buyer-utilitarian welfare functions with decreasing Pareto weights. Under such welfare functions, the value to the designer of an additional unit of surplus is lower for high-types than for low-types. Decreasing Pareto weights can endogenously arise under a utilitarian objective if buyers with a low willingness to pay have a higher marginal utility of money than those with a high willingness to pay (Condorelli, 2013; Dworczak et al., 2021), or if buyers are identified by unobserved welfare weights that negatively correlate with their willingness to pay (Akbarpour ⓡ al., 2024b).
@@
belongs to $\mathcal{R}$ if it is concave, and belongs to $\overline{\mathcal{R}}$ if it is strictly concave. Increasing concave transformations of surplus correspond to the class of welfare functions usually considered in optimal taxation (Mirrlees, 1971) and insurance (Rothschild and Stiglitz, 1976). [^17] Garcia and Tsur (2021) characterize optimal segmentations of a competitive insurance market. Such increasing concave transformations of surplus thus arise endogenously under a utilitarian objective. Concavity reflects equity concerns, as it implies diminishing marginal returns to welfare from additional units of surplus. Any designer whose welfare function is given by Equation 2 for some increasing concave $u$ would therefore prefer high prices to be charged to high-types than to low-types.
Note also that any welfare function which is a combination of Examples 1 and 2, given by
@@
For any $w\in\mathcal{R}$ and $\mu\in\Delta(\Theta)$, we say that $\sigma$ is $\mathrm{P}_{w,\mu}$-optimal if it solves the optimization program $\mathrm{P}_{w,\mu}$. Furthermore, we say that a segmentation is redistributive if there exist $w\in\mathcal{R}$ and $\mu\in\Delta(\Theta)$ such that $\sigma$ is $\mathrm{P}_{w,\mu}$-optimal.
3 Redistributive segmentations and welfare
3.1 Efficiency
We start by proving that it is without loss of generality to search for optimal segmentations among the efficient ones. A segmentation is efficient if it leads the seller to serve all the buyers. We define $\Omega=\bigl{\{}(\theta,p)\in\Theta\times P\ |\ \theta\geq p\bigr{\}}$ as the set of market outcomes where buyers consume. Hence, a segmentation is efficient if it is supported on $\Omega$. For any $\mu\in\Delta(\Theta)$, we let
$\Sigma_{\Omega}(\mu)\coloneqq\bigl{\{}\sigma\in\Sigma(\mu)\ |\ \operatorname{supp}(\sigma)\subseteq\Omega\bigr{\}},$
be the set of efficient segmentations of $\mu$.
@@ 3.1 Efficiency
For any $w\in\mathcal{R}$ and $\mu\in\Delta(\Theta)$, there exists a $\mathrm{P}_{w,\mu}$-optimal segmentation that belongs to $\Sigma_{\Omega}(\mu)$.
@@
Let $\mu\in\Delta(\Theta)$ and consider $\sigma\in\Sigma(\mu)$ such that $\operatorname{supp}(\sigma)\not\subset\Omega$. There must therefore exist $(\theta,p)\in\operatorname{supp}(\sigma)$ such that $\theta<p$. Let us contemplate transferring the entire mass of buyers at $(\theta,p)$ to $(\theta,\theta)$. This transfer violates neither $\mathrm{C}_{\mu}$ nor Ob. Indeed, the total mass of type $\theta$ buyers still equals $\mu(\theta)$ so that condition $\mathrm{C}_{\mu}$ is preserved. Furthermore, this buyer transfer relaxes the obedience constraints in both segments $\theta$ and $p$. Finally, since $w(\theta,p)=0\leq w(\theta,\theta)$, the transfer must increase the aggregate welfare. In fact, aggregate welfare weakly increases without hurting the seller or the buyers—a (weak) Pareto improvement. ∎
@@
Lemma 1 establishes that segmenting a market in a redistributive way never hinders efficiency. The intuition behind that result is similar to Haghpanah and Siegel’s (2023) main result. Namely, any inefficient segmentation is Pareto-dominated by some efficient one, which improves aggregate welfare as redistributive welfare functions are aligned with consumer surplus.
Operationally, Lemma 1 implies that there is no loss of generality in considering only efficient segmentations as candidates for $\mathrm{P}_{w,\mu}$-optimality. We can therefore restate the segmentation design problem $\mathrm{P}_{w,\mu}$ as choosing a welfare-maximizing efficient segmentation:
$\max_{\sigma\in\Sigma_{\Omega}(\mu)}\sum_{(\theta,p)\in\Omega}w(\theta,p)\,\sigma(\theta,p)$ ($\mathrm{P}^{\star}_{w,\mu}$)
3.2 The redistributive order
We propose an ordering of efficient segmentations based on their redistributivity, which we call the redistributive order. This ordering will help characterize qualitative properties of redistributive segmentations in Sections 3.3 and 3.4.
Definition 1.
Let $\mu\in\Delta(\Theta)$ and $\sigma,\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$. We say that $\sigma$ is more redistributive than $\sigma^{\prime}$, denoted $\sigma\succsim_{\mathcal{R}}\sigma^{\prime}$, if
$\forall w\in\mathcal{R},\quad\sum_{(\theta,p)\in\Omega}w(\theta,p)\,\sigma(\theta,p)\geq\sum_{(\theta,p)\in\Omega}w(\theta,p)\,\sigma^{\prime}(\theta,p),$
We call $\succsim_{\mathcal{R}}$ the redistributive order over $\Sigma_{\Omega}(\mu)$ and denote as $\succ_{\mathcal{R}}$ its asymmetric part. [^18] One can show that $\succsim_{\mathcal{R}}$ is a partial order, that is, a transitive, reflexive and antisymmetric binary relation. While the first two conditions are easy to verify from the definition, the fact that it is antisymmetric follows from Theorem 1 below.
Two efficient segmentations are ranked according to the redistributive order if one leads to greater aggregate welfare for any redistributive welfare function. Given any market $\mu$, a more redistributive segmentation of $\mu$ is thus precisely one that is preferred ex-ante by any designer with a redistributive objective. Note that $\succsim_{\mathcal{R}}$ is not a complete order. One specific segmentation may produce higher aggregate welfare than another for certain redistributive welfare functions, while the latter produces higher aggregate welfare than the former for different redistributive welfare functions.
The redistributive order is defined in the same spirit as the (bivariate) supermodular stochastic order (Shaked and Shanthikumar, 2007, Section 9.A.4.), with two notable differences. First, in addition to being supermodular, redistributive welfare functions are also decreasing in their second dimension (corresponding here to the price/segment). Second, efficient segmentations correspond to bivariate distributions such that only the first marginal is prescribed. We leverage this connection by characterizing the redistributive order in a similar way Müller (2013) and Meyer and Strulovici (2015) characterize the supermodular stochastic order. Namely, we characterize $\succsim_{\mathcal{R}}$ in terms of mass transfers.
A buyer mass transfer (in short a transfer) is a function $t\colon\Omega\to\mathbb{R}$ describing buyers’ mass movements within $\Omega$. Among all the possible transfers, the following two classes will enable us to characterize $\succsim_{\mathcal{R}}$.
Definition 2 (Downward buyer transfer).
A buyer transfer $t\colon\Omega\to\mathbb{R}$ is downward if there exists $\delta\geq 0$ and $\tilde{\theta},p^{\prime},p^{\prime\prime}\in\Theta$ such that $p^{\prime}<p^{\prime\prime}$, and
Definition 3 (Redistributive buyer transfer).
A buyer transfer $t\colon\Omega\to\mathbb{R}$ is redistributive if there exists $\varepsilon\geq 0$ and $(\theta^{\prime},p^{\prime\prime}),(\theta^{\prime\prime},p^{\prime})\in\Omega$ such that $\theta^{\prime}<\theta^{\prime\prime}$, $p^{\prime}<p^{\prime\prime}$, and
Figure 2 illustrates the downward and redistributive transfer operations on an arbitrary segmentation.
$\theta$$p$$\delta$$\varepsilon$$\varepsilon$
@@ 3.2 The redistributive order
Downward transfers carry a certain mass of a given type of consumer from one segment to another with a lower price. They are always desirable since the redistributive welfare functions decrease in price. Redistributive transfers exchange a mass of some type assigned to a particular segment for the same mass of higher types assigned to a segment with a lower price. These transfers are also always desirable since condition R ensures that the welfare gain from lowering prices for low-types outweighs the welfare loss from increasing prices for high-types.
Therefore, any combination of downward and redistributive transfers increases aggregate welfare if the welfare function is redistributive. As our first result shows, these combinations of transfers are also the only ones to increase aggregate welfare for any redistributive objective. Let $\mathcal{T}$ be the set of all downward and redistributive buyer transfers.
Theorem 1 (Redistributive order).
For any $\mu\in\Delta(\Theta)$ and $\sigma,\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$, $\sigma\succsim_{\mathcal{R}}\sigma^{\prime}$ if and only if there exists $n\in\mathbb{N}$, $n\neq 0$, and $(t_{i})_{i\in\{1,\dots,n\}}\in\mathcal{T}^{n}$ such that
$\sigma=\sigma^{\prime}+\sum_{i=1}^{n}t_{i}.$
Theorem 1 states that one segmentation is less redistributive than another if and only if one can move from the former to the latter by performing a finite number of the elementary transfers described in Definitions 2 and 3. Its proof, which can be found in Appendix A, relies on duality arguments and follows a similar approach to the proofs of Theorem 2.4.1 in Müller (2013) and Theorem 1 in Meyer and Strulovici (2015).
3.3 Order-maximal segmentations and optimality
We now leverage Theorem 1 to uncover properties of redistributive segmentations. For any $\mu\in\Delta(\Theta)$, we say that a segmentation $\sigma\in\Sigma_{\Omega}(\mu)$ is $\succsim_{\mathcal{R}}$-maximal if there exists no other segmentation $\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$ such that $\sigma^{\prime}\succ_{\mathcal{R}}\sigma$.
Proposition 1 (Optimality).
For any $\mu\in\Delta(\Theta)$, the following two statements are true.
  1. (i)
    For any $w\in\mathcal{R}$, there exists a $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation of $\mu$ that is $\succsim_{\mathcal{R}}$-maximal. Furthermore, if $w\in\overline{\mathcal{R}}$, then any $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation of $\mu$ is $\succsim_{\mathcal{R}}$-maximal.
  2. (ii)
    Conversely, if $\sigma\in\Sigma_{\Omega}(\mu)$ is $\succsim_{\mathcal{R}}$-maximal, then there exists $w\in\mathcal{R}$ such that $\sigma$ is $\mathrm{P}^{\star}_{w,\mu}$-optimal.
Proposition 1 establishes a correspondence between optimal segmentations and the maximal elements of the redistributive order. Claim (i) asserts that it is without loss of generality to restrict attention to $\succsim_{\mathcal{R}}$-maximal elements when looking for redistributive segmentations (and even necessary to do so when the welfare function is strictly redistributive). This implication is a direct consequence of Theorem 1, since for any non-maximal segmentation there exists at least one sequence of transfers in $\mathcal{T}$ which increases aggregate welfare for any redistributive welfare function without violating any of the constraints defining $\Sigma_{\Omega}(\mu)$. Claim (ii) states that for any aggregate market, a $\succsim_{\mathcal{R}}$-maximal element can be rationalized as an optimal segmentation for some redistributive welfare function. Its proof can be found in Appendix B and relies on a separation argument.
Building on Proposition 1, we show that the saturation of a particular subset of the seller’s obedience constraints characterizes redistributive segmentations. [^19] Haghpanah and Siegel (2024) use a similar notion to our downward transfers to characterize when a segmentation is stable (meaning that no coalition of buyers can obtain a better outcome by deviating to another segment). In particular, they define a segmentation as saturated if transferring any positive mass of buyers to a given segment from a segment with a higher price increases the price in the lower segment. They show that a segmentation is stable if and only if it is efficient and saturated.
Proposition 2 (Characterization).
For any $\mu\in\Delta(\Theta)$, a segmentation $\sigma\in\Sigma_{\Omega}(\mu)$ is $\succsim_{\mathcal{R}}$-maximal if and only if the following conditions hold:
  1. (i)
    At any segment $p\in\operatorname{supp}(\sigma_{P})$ such that $p<\max\operatorname{supp}(\sigma_{P})$, the seller is indifferent between charging $p$ and some other price $q>p$;
  2. (ii)
    For any segment $p\in\operatorname{supp}(\sigma_{P})$ and any $\theta\geq p$, if $\theta\in\operatorname{supp}\bigl{(}\sigma(\,\cdot\,|\,p)\bigr{)}$ then, at any segment $p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p<p^{\prime}\leq\theta$, the seller is indifferent between charging $p^{\prime}$ and $\theta$.
The proof of Proposition 2 can be found in Appendix C. Let us explain its logic. From Theorem 1 we know that, for any $\mu\in\Delta(\Theta)$, a $\succsim_{\mathcal{R}}$-maximal segmentation $\sigma\in\Sigma_{\Omega}(\mu)$ is precisely one on which performing any sequence $(t_{i})_{i\in\{1,\dots,n\}}\in\mathcal{T}^{n}$ of downward or redistributive transfers would imply that $(\sigma+\sum_{i=1}^{n}t_{i}\bigl{)}$ violates Ob. [^20] We can restrict attention to how downward and redistributive transfers impact the seller’s obedience constraints since, for any market $\mu\in\Delta(\Theta)$, segmentation $\sigma\in\Sigma_{\Omega}(\mu)$ and transfer $t\in\mathcal{T}$ sufficiently small, the resulting distribution $\sigma+t$ must satisfy $\mathrm{C}_{\mu}$. Indeed, the transfers belonging to $\mathcal{T}$ preserve the total mass of each buyer type, as $\sum_{p\in P}t(\theta,p)=0$ for any $\theta\in\Theta$. We start by establishing in Lemma 6 that, for a segmentation to be $\succsim_{\mathcal{R}}$-maximal it is sufficient that any single transfer $t\in\mathcal{T}$ induces $\sigma+t$ to violate Ob. For any $\mu\in\Delta(\Theta)$ and $\sigma\in\Sigma_{\Omega}(\mu)$, we thus say that a transfer $t$ is feasible for $\sigma$ if it preserves Ob, that is, if $(\sigma+t)\in\Sigma_{\Omega}(\mu)$. We conversely say that a transfer $t\in\mathcal{T}$ is unfeasible for $\sigma$ if it is not feasible.
@@ 3.3 Order-maximal segmentations and optimality
Let us investigate the conditions under which a particular downward transfer is unfeasible for some segmentation. Let $\mu\in\Delta(\Theta)$ and $\sigma\in\Sigma_{\Omega}(\mu)$ be such that there exist two segments $p^{\prime},p^{\prime\prime}\in\operatorname{supp}(\sigma_{P})$ with $p^{\prime\prime}>p^{\prime}$, and let $\tilde{\theta}\in\operatorname{supp}\bigl{(}\sigma(\,\cdot\,|\,p^{\prime\prime})\bigr{)}$. Consider transferring downward a mass $\delta$ of types $\tilde{\theta}$ from the segment $p^{\prime\prime}$ to $p^{\prime}$, as illustrated in Figure 3(a).

### v5

@@ 2. Model
The value $w(\theta,p)$ thus represents the individual contribution to the aggregate welfare when a buyer of type $\theta$ is assigned to a segment where the seller charges the price $p$.
We restrict attention to welfare functions satisfying three conditions. First, the value from assigning a buyer to a segment where he does not consume is zero, and the value from assigning a buyer to a segment where he consumes is non-negative. That is, $w(\theta,p)=0$ if $\theta<p$, and $w(\theta,p)\geq 0$ if $\theta\geq p$. Second, assigning any buyer to a segment where the seller charges a lower price always improves welfare. That is, $p\mapsto w(\theta_{k},p)$ is a non-increasing function on $\{\theta_{1},\dots,\theta_{k}\}$ for all $k\in\{1,\dots,K\}$. The two preceding conditions capture the fact that social welfare aligns with consumer surplus. [^20] It is straightforward to verify that the buyer-utilitarian welfare function, given by $w(\theta,p)=(\theta-p)\mathds{1}_{\theta\geq p}$ for all $(\theta,p)\in\Theta\times P$, satisfies these two minimal conditions. The third condition captures the designer’s redistributive concerns. We require that any welfare function be supermodular conditional on buyers consuming. That is, for any $(\theta,p),(\theta^{\prime},p^{\prime})\in\Theta\times P$ such that $\theta^{\prime}>\theta\geq p^{\prime}>p$,
@@
Condition R ensures that any price discount from $p^{\prime}$ to $p$ increases welfare more when it is granted to a low-type $\theta$ rather than to a high-type $\theta^{\prime}$.
We say that a welfare function is redistributive if it satisfies the three conditions above, and denote by $\mathcal{R}$ the class of those welfare functions. Furthermore, we say that $w$ is strictly redistributive if it satisfies the condition R strictly on all $\Theta\times P$, and is such that $p\mapsto w(\theta_{k},p)$ is strictly decreasing on $\{\theta_{1},\dots,\theta_{k}\}$ for all $k\in\{1,\dots,K\}$. We denote the class of strictly redistributive welfare functions by $\overline{\mathcal{R}}$.
In Examples˜1 and 2 below, we highlight that the class of redistributive welfare functions encompasses two common specifications.
@@
belongs to $\mathcal{R}$ whenever $\theta\mapsto\lambda(\theta)$ is non-increasing, and belongs to $\overline{\mathcal{R}}$ whenever $\theta\mapsto\lambda(\theta)$ is strictly decreasing. The class of welfare functions given by 1 for some decreasing function $\lambda$ corresponds to the family of weighted buyer-utilitarian welfare functions with decreasing Pareto weights. Under such welfare functions, the value to the designer of an additional unit of surplus is lower for high-types than for low-types. Decreasing Pareto weights can endogenously arise under a utilitarian objective if buyers with a low willingness to pay have a higher marginal utility of money than those with a high willingness to pay (Condorelli2013; dworczak21), or if buyers are identified by unobserved welfare weights that negatively correlate with their willingness to pay (akbarpour2024redistributive).
@@
belongs to $\mathcal{R}$ if $u$ is concave, and belongs to $\overline{\mathcal{R}}$ if $u$ is strictly concave. Increasing concave transformations of surplus correspond to the class of welfare functions usually considered in optimal taxation (mirrlees1971exploration) and insurance (Rothschild1976). Concavity reflects equity concerns, as it implies diminishing marginal returns to welfare from additional units of surplus.
Note also that any welfare function which is a combination of Examples˜1 and 2, given by
@@
3. Redistributive Segmentations and Welfare
We start by proving a preliminary result that simplifies the analysis: it is without loss of generality to restrict the search for optimal redistributive segmentations to the efficient ones. We then characterize the optimal redistributive segmentations. Finally, we discuss the economic implications of our characterization. First, we show that optimal redistributive segmentations must exhibit a monotonic structure that induces the seller to price progressively. Second, we show that when the redistributive concerns are strong, optimal redistributive segmentations can lead the seller to obtain a strictly positive share of the surplus.
3.1.  Efficiency
A segmentation is efficient if it leads the seller to serve all the buyers. Let $\Omega=\bigl\{(\theta,p)\in\Theta\times P\ |\ \theta\geq p\bigr\}$. For any $\mu\in\Delta(\Theta)$, the set of efficient segmentations of $\mu$ is defined by
$\Sigma^{\mathrm{eff}}(\mu)=\bigl\{\sigma\in\Sigma(\mu)\ |\ \operatorname{supp}(\sigma)\subseteq\Omega\bigr\}.$
@@ 3.1.  Efficiency
For any $w\in\mathcal{R}$ and $\mu\in\Delta(\Theta)$, there exists an optimal segmentation for the problem $\mathrm{P}_{w,\mu}$ that belongs to $\Sigma^{\mathrm{eff}}(\mu)$.
@@
Let $\mu\in\Delta(\Theta)$ and consider $\sigma\in\Sigma(\mu)$ with $\operatorname{supp}(\sigma)\not\subset\Omega$. Therefore, there must exist $(\theta,p)\in\operatorname{supp}(\sigma)$ such that $\theta<p$. Consider reallocating the entire mass of buyers $\sigma(\theta,p)$ to $(\theta,\theta)$. This reallocation violates neither $\mathrm{C}_{\mu}$ nor Ob. The total mass of buyers of type $\theta$ is still equal to $\mu(\theta)$ so $\mathrm{C}_{\mu}$ is preserved. Furthermore, this reallocation relaxes Ob in both segments $\theta$ and $p$. Finally, since $w(\theta,p)=0\leq w(\theta,\theta)$, the aggregate welfare must increase. In fact, aggregate welfare weakly increases without hurting the seller or the buyers—a (weak) Pareto improvement. ∎
@@
Lemma˜1 establishes that segmenting a market in a redistributive way never hinders efficiency. The intuition behind that result is similar to hagpanah_siegel_jpe’s (hagpanah_siegel_jpe) main result: any inefficient segmentation is Pareto-dominated (with respect to the seller’s profit and consumer surplus) by some efficient one. For our class of welfare functions, such Pareto improvements always increase aggregate welfare, since $w(\theta,\,\cdot\,)$ is a decreasing function for each $\theta\in\Theta$.
Operationally, Lemma˜1 implies that we can restate the segmentation design problem $\mathrm{P}_{w,\mu}$ as follows:
$\max_{\sigma\in\Sigma^{\mathrm{eff}}(\mu)}\;\sum_{(\theta,p)\in\Theta\times P}w(\theta,p)\,\sigma(\theta,p).$ ($\mathrm{P}^{\star}_{w,\mu}$)
For any $w\in\mathcal{R}$ and $\mu\in\Delta(\Theta)$, we say that $\sigma$ is $\mathrm{P}^{\star}_{w,\mu}$-optimal if it solves the optimization program $\mathrm{P}^{\star}_{w,\mu}$. We call optimal redistributive segmentations all segmentations that are $\mathrm{P}^{\star}_{w,\mu}$-optimal for some $\mu\in\Delta(\Theta)$ and $w\in\mathcal{R}$.
3.2.  Characterization of Optimal Segmentations
The characterization of $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentations, stated below as Theorem˜1, is given in terms of the conditions described in Definition˜1 (see below for a detailed explanation).
Definition 1 (Redistributive Saturation).
Let $\mu\in\Delta(\Theta)$. A segmentation $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$ is redistributively saturated (saturated for short) if the following conditions hold:
  1. (a)
    In any segment $p\in\operatorname{supp}(\sigma_{P})$ such that $p<\max\operatorname{supp}(\sigma_{P})$, the seller is indifferent between charging $p$ and some other price $q>p$;
  2. (b)
    Let $p\in\operatorname{supp}(\sigma_{P})$ and $\theta\in\operatorname{supp}\bigl(\sigma(\,\cdot\,|\,p)\bigr)$. Then in every segment $p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p<p^{\prime}\leq\theta$ the seller is indifferent between charging $p^{\prime}$ and $\theta$.
An efficient segmentation is saturated if (a) in every segment, except maybe for the one with the highest price, the seller faces more than one optimal price, and (b) if some consumers of type $\theta$ belong to a segment with price $p$, then consumers of type $\theta$ must be sufficiently represented in higher segments so that $\theta$ is an optimal price in those segments.
Theorem 1 (Optimal redistributive segmentations).
For any $\mu\in\Delta(\Theta)$, the two following claims are true:
  1. (i)
    If $w\in\mathcal{R}$, then there exists a $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation which is saturated. Furthermore, if $w\in\overline{\mathcal{R}}$, then every $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation is saturated.
  2. (ii)
    Conversely, if $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$ is saturated, then there exists $w\in\mathcal{R}$ such that $\sigma$ is $\mathrm{P}^{\star}_{w,\mu}$-optimal.
The above result establishes that the set of saturated segmentations and optimally redistributive ones are (almost) equal. Claim (i) states that an optimal segmentation must be saturated if the welfare function is strictly redistributive ($w\in\overline{\mathcal{R}}$). If $w\in\mathcal{R}$ is not strictly redistributive, then there may be indifferences that make some non-saturated segmentations also optimal, but a saturated optimal one always exists. Claim (ii) states that, conversely, any saturated segmentation is optimal for some redistributive objective. As a result, any property that is satisfied by saturated segmentations is also satisfied by optimal redistributive segmentations. We leverage this result to draw economic implications of optimal redistributive segmentations in Sections˜3.3 and 3.4; Section˜3.2.2 outlines its proof. [^21] From now on, with slight abuse of terminology, we use the terms optimally redistributive segmentation and saturated segmentation interchangeably.
3.2.1. Intuitive Explanation of Theorem˜1
Theorem˜1 is best understood in light of what saturated segmentations achieve. Saturated segmentations are those for which no reallocation of buyers between segments could increase the expectation of every redistributive welfare function without leading to a violation of the seller’s obedience constraints. At this stage, two questions are in order: (1) Which type of reallocation increases the expectation of every redistributive welfare function? (2) When is it not possible to perform such reallocations without violating the seller’s obedience constraints?
Proposition˜1, formally stated in Section˜3.2.2, answers question (1) by identifying the two marginal reallocations of buyers that increase any redistributive welfare function. First, downward buyer transfers (Definition˜3) move buyers to segments with lower prices. Second, redistributive buyer transfers (Definition˜4) swap a given mass of low-type buyers in a high-price segment for an equal mass of higher-type buyers in a lower-price segment. Figure˜2 illustrates both transfers on an efficient segmentation.
[…]
@@ 3.2.1. Intuitive Explanation of Theorem˜1
Downward transfers are desirable under redistributive welfare functions as they decrease the price charged to some buyers. Redistributive transfers are also desirable under redistributive welfare functions due to condition R. The main contribution of Proposition˜1 is to establish the converse: any buyer reallocation that increases the expected value of all redistributive welfare functions must necessarily consist of some (finite) combination of downward and redistributive buyer transfers. This result provides a complete characterization of welfare-improving buyer reallocations under redistributive objectives.
While downward and redistributive buyer transfers are the key to making segmentations more redistributive, these transfers interfere with the seller’s obedience constraints. Reallocating buyers through downward or redistributive transfers can lead the seller to deviate to a higher price in one or more segments. Lemmas˜2 and 3, also formally stated in Section˜3.2.2, answer question (2). They show that saturation is precisely the condition that ensures that no downward or redistributive transfer (or a combination thereof) can be performed without causing the seller to deviate from the recommended prices.
@@
To get a sense of why, let us first investigate the conditions under which a particular downward transfer is infeasible for some segmentation. Consider transferring some mass of buyers of type $\tilde{\theta}$ from a higher-price segment $p^{\prime\prime}$ to a lower-price segment $p^{\prime}$, as illustrated in Figure˜3(a).